Markus Company’s common stock sold for $3.75 per share at the end of this year. The company paid a common stock dividend of $0.60 per share this year. It also provided the following data excerpts from this year’s financial statements:     

Ending
Balance
Beginning
Balance
Cash $ 41,000 $ 36,200
Accounts receivable $ 74,000 $ 58,700
Inventory $ 64,900 $ 74,000
Current assets $ 179,900 $ 168,900
Total assets $ 590,000 $ 624,000
Current liabilities $ 70,500 $ 63,000
Total liabilities $ 160,000 $ 144,000
Common stock, $1 par value $ 139,000 $ 139,000
Total stockholders’ equity $ 430,000 $ 480,000
Total liabilities and stockholders’ equity $ 590,000 $ 624,000

     

This Year
Sales (all on account) $ 865,000
Cost of goods sold $ 501,700
Gross margin $ 363,300
Net operating income $ 206,250
Interest expense $ 11,000
Net income $ 136,675

11. What is the company’s operating cycle?
Answer #1

A Company's Operating Cycle is the Sum of the following:-

1. The Days In Inventory (365 Days/Inventory Turnover Ratio)

2. The Average Collection Period (365 Days/Accounts Receivable Turnover Ratio)

So firstly we need to Calculate Inventory Turnover Ratio and Days In Inventory and Accounts Receivable Turnover Ratio and Average Collection Period.

Computation of Inventory Turnover Ratio:-

Inventory Turnover Ratio calculated by dividing Cost of Goods Sold by Average Inventory.

Average Inventory=(Beginning Inventory + Ending Inventory)/2

=$(74,000+64,900)/2

=$(138,900/2)

=$69,450

Inventory Turnover Ratio=(Cost of Goods Sold /Average Inventory)

=($501,700/69,450)

=7.22 Times

Computation of Days In Inventory:-

Days In Inventory=(365 Days/Inventory Turnover Ratio)

=(365/7.22) Days

=50.55 Days

Computation of Accounts Receivable Turnover Ratio:-

Accounts Receivable Turnover Ratio calculated by dividing Net Credit Sales by Average Accounts Receivable.

Average Accounts Receivable=( Beginning Accounts Receivable+ Ending Accounts Receivable)/2

=$(58,700+74,000)/2

=$(132,700/2)

=$66,350

Accounts Receivable Turnover Ratio=( Net Credit Sales/ Average Accounts Receivable)

=($865,000/66,350)

=13.04 Times

Computation of Average Collection Period:-

Average Collection Period=(365 Days/ Accounts Receivable Turnover Ratio)

=(365/13.04) Days

=27.99 Days

Operating Cycle of Company:-

Operating Cycle=(Days In Inventory+Average Collection Period)

=(50.55+27.99) Days

=78.54 Days

So the Operating Cycle of Company is 78.54 Days or 79 Days.